Consider three charges ๐‘ž1=1.5 ๐œ‡๐ถ, ๐‘ž2ย =5๐œ‡๐ถ, and ๐‘ž3=-3 ๐œ‡๐ถ located at (๐‘ฅ1,๐‘ฆ1,๐‘ง1)=(1,โˆ’2,4) ๐‘š, (๐‘ฅ2,๐‘ฆ2,๐‘ง2)=(2,-1, 3) ๐‘š, andย (๐‘ฅ3,๐‘ฆ3,๐‘ง3)=(3, 3, -4) ๐‘š, respectively. Find the total net forceย F

โƒ—ย 

1

Fโ†’1ย on charge ๐‘ž1ย due to electric fieldsย E

โƒ—ย 

2

Eโ†’2ย (induce by ๐‘ž2)ย andย E

โƒ—ย 

3

Eโ†’3ย (induce by ๐‘ž3) . Note that by superposition of vectors:


1
Expert's answer
2021-03-21T11:31:05-0400

Given,

"q_1=1.5\mu C"

"q_2=5\mu C"

"q_3=-3\mu C"

(๐‘ฅ1,๐‘ฆ1,๐‘ง1)=(1,โˆ’2,4) ๐‘š,

"r_1=\hat{i}-2\hat{j}+4\hat{k}"

(๐‘ฅ2,๐‘ฆ2,๐‘ง2)=(2,-1, 3) ๐‘š,

"r_2=2\hat{i}-\hat{j}+3\hat{k}"


Andย (๐‘ฅ3,๐‘ฆ3,๐‘ง3)=(3, 3, -4) ๐‘š

"r_3=3\hat{i}+3\hat{j}-4\hat{k}"

"d_{12}=\sqrt{(2-1)^2+(-1+2)^2+(3-4)^2}=\sqrt{1+1+1}=\sqrt{3}"

"\overrightarrow{d_{12}}=\hat{i}+\hat{j}-\hat{k}"


"d_{13}=\sqrt{(3-1)^2+(3+2)^2+(-4-4)^2}=\sqrt{4+25+64}=\sqrt{93}"

"\overrightarrow{d_{13}}=2\hat{i}+5\hat{j}-8\hat{k}"


"d_{23}=\sqrt{1+4^2+8^2}=\sqrt{81}=9"

"\overrightarrow{d_{23}}=\hat{i}+4\hat{j}-8\hat{k}"

Hence, the required force "F_{12}=\frac{9\times 10^{9}\times 1.5\times 10^{-6}\times 5\times 10^{-6}}{3}\frac{\hat{i}+\hat{j}-\hat{k}}{\sqrt{3}}" N

"=7.5\sqrt{3}\times 10^{-3}(\hat{i}+\hat{j}-\hat{k})N"


"F_{13}=-\frac{9\times 10^{9}\times 1.5\times 3\times 10^{-12}}{93}\frac{2\hat{i}+5\hat{j}-8\hat{k}}{\sqrt{93}}"

"=-0.044\times 10^{-3}(2\hat{i}+5\hat{j}-8\hat{k})"

"F_{net}=F_{12}+F_{13}=7.5\sqrt{3}\times 10^{-3}(\hat{i}+\hat{j}-\hat{k})-0.044\times 10^{-3}(2\hat{i}+5\hat{j}-8\hat{k})"

"=12.9\hat{i}+12.77\hat{j}-12.63\hat{k}"

"F_{23}=\frac{9\times 15\times 10{-3}}{81}\frac{\hat{i}+4\hat{j}-8\hat{k}}{9}"